A  143kg object is dropped from a height of 3m above a spring with a constant 11545N/m. How much (in cm) will the spring be compressed?



1
Expert's answer
2021-04-26T17:12:58-0400

According to the energy conservation law, the initial gravitational potential energy of the object was converted into the elastic potential energy of the spring:


"E_p = E_e"

The gravitational potential energy is:


"E_p = mg(h + x)"

where "m = 13kg" is the mass of the object, "g = 9.8 m\/s^2" is the gravitational acceleration "h = 3m" is the height above the spring, and "x" is the spring compression. Here the fully compressed spring is taken as the zero level for gravitational potential energy.

The elastic potential energy is:


"E_e =\dfrac{kx^2}{2}"

where "k = 11545N\/m" is the spring constant.

Thus, obtain the following equation:


"mg(h+x) = \dfrac{kx^2}{2}\\
\dfrac{kx^2}{2} - mgx - mgh = 0"

Solving the quadratic equation, obtain:


"x \approx 0.27\space \text{or} \space x \approx -0.25"

Since we are interested in comression, take only positive value.


Answer. 0.27 m.